OpenStudy (anonymous):

HELP!!!!! If a freely falling rock were equipped with a speedometer, by how much would its speed readings increase with each second IF it were on a planet where g = 20 m/s2?

OpenStudy (aaronq):

 \(\large v_f=v_i+a\Delta t\) \(a= 20 ~m/s^2\) initial velocity time final velocity \(v_i=0\) t=1 sec \(v_f=0+(20~m/s^2)(1s)=20~m/s\) \(v_i=0\) t=2 sec \(v_f=0+(20~m/s^2)(2s)=40~m/s\) \(v_i=0\) t=3 sec \(v_f=0+(20~m/s^2)(3s)=60~m/s\) \(v_i=0\) t=4 sec \(v_f=0+(20~m/s^2)(4s)=80~m/s\)

OpenStudy (anonymous):

@aaronq - I have what you wrote but I still can't seem to come to a conclusion =/

OpenStudy (aaronq):

notice that each second in freefall increases the speed by 20 m/s

OpenStudy (aaronq):

from t=1 s to t=2 s, there is a net 20 m/s increase

OpenStudy (anonymous):

Ahhhh makes sense. Thank you sooooo much!Aa

OpenStudy (anonymous):

@aaronq You rock!

OpenStudy (aaronq):

no problem! glad it helped